OpenStudy (anonymous):

square root -343 express in terms of i (the "i" is italicized)

OpenStudy (anonymous):

\[\sqrt{-343} = \sqrt{343(-1)} = \sqrt{343}\sqrt{-1} = \sqrt{343}i\]

OpenStudy (anonymous):

Which can be simplified to \(7 \sqrt{7}i\) as 343 = 7^3.